OpenStudy (anonymous):

w = 3, x = 6, y = 5, and z = 9 y2-5z

OpenStudy (anonymous):

5(2)-5(9) 10-54 -44

jimthompson5910 (jim_thompson5910):

\(\Large y^2 - 5z\) \(\Large (5)^2 - 5(9)\) ... replace y with 5 and z with 9 Then evaluate

OpenStudy (anonymous):

i don't understand what is the need of ( w=3 & x=6) !!!

jimthompson5910 (jim_thompson5910):

probably for a previous (or next) problem

OpenStudy (anonymous):

oh it's y squared? ok then

OpenStudy (anonymous):

replace "y" with 5 and "z" with 9 and do the math

OpenStudy (anonymous):

(5)^2-5(9)=25-45=20